OpenStudy (anonymous):

If 2x^2+10x+6=6, which of the following could be the value of x+5? I. 1 II. 5 III.10 The Answer is 5 but how is that the answer?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 2x^2+10x+6=6\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \implies 2x^2+10x=0\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Alright subtract six. Would you then solve for x?

OpenStudy (mathmath333):

yes of cos

OpenStudy (anonymous):

Of cosine?

OpenStudy (mathmath333):

lol , i mean of course

OpenStudy (anonymous):

Oh my bad hah. Alright so

OpenStudy (anonymous):

\[2x(x+5)\]

OpenStudy (mathmath333):

sry this one \(\large\tt \begin{align} \color{black}{ \implies 2x(x+5)=0\hspace{.33em}\\~\\ \implies x(x+5)=0\hspace{.33em}\\~\\ \implies x=0~~~or~~~x+5=0\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Alright thta's what I got x=0 and x=-5

OpenStudy (mathmath333):

yes but the question asks \(\large\tt \begin{align} \color{black}{ (x+5)=?\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

oh yes that's right

OpenStudy (anonymous):

That's where I'm confused.

OpenStudy (mathmath333):

we found \(\large\tt \begin{align} \color{black}{ x=0\hspace{.33em}\\~\\ }\end{align}\) add \(5\) to both sides

OpenStudy (anonymous):

so x+5=5?

OpenStudy (anonymous):

oh dam haha that's the answer XD

OpenStudy (mathmath333):

yes that should be it the options dont have \(0\) as the answer

OpenStudy (anonymous):

Thank you very much!

OpenStudy (mathmath333):

yw

OpenStudy (mathmath333):

is is probably a trick question

OpenStudy (mathmath333):

*this is